# Expanding Chatbot Knowledge in Customer Service: Context-Aware Similar Question Generation Using Large Language Models

### v2

@@ 4.1 Main Results
The results for generating 20 questions are presented in Table 2. Most methods demonstrate high precision scores, where the generated questions maintain strong semantic consistency with the source question. In contrast, baseline methods perform poorly in terms of recall, suggesting that the generated questions are largely constrained within a narrow semantic space around the source question. Methods relying on in-context learning exhibit poor performance in both precision and recall due to their limited knowledge of the specific SQG task. By fine-tuning the LLM with a naive one-to-one objective (ChatGLM2-FT), the semantic diversity is enhanced without distorting precision, which emphasizes the significance of task-specific adaption. Integrating the proposed one-to-many training objective leads to a higher recall score of 0.8377, while the Intention-Enhanced method presented superior but limited improvement in these quantitative metrics.
With human evaluation, we show that general-purpose task-generation models SimBERT and RoFormer-Sim exhibit poor performance, with only 20% of the generated questions meeting the acceptance criteria. ChatGLM2-FT increases this proportion to 37.9%, but still remains largely redundant and does not fulfill practical application needs. While the Context-Aware method excels in quantitative metrics, it only shows a modest increase in the acceptance ratio. Introducing the customer’s intention significantly broadens the exploration space of generated questions. This underscores the vital role of contextual information in moving beyond questions uniformly centered around the source question while maintaining high consistency with both the source question and the intended answer.
Lastly, the character-level diversity indicates that ChatGLM2-FT and the proposed batch generation methods achieve higher Distinct scores than SimBERT, RoFormer-Sim, and ChatGLM2 (Few-Shot). Although the zero-shot setting achieves the highest Distinct score, it comes at the cost of significantly lower consistency due to its unconditioned generation process, which is not a preferred behavior.
Figure 3: Performance comparison of different methods when generating a varying number of questions.
@@
In real applications, the desired quantity of similar questions often falls within the range of several tens to hundreds. Therefore, we systematically examine the performance of the proposed method by generating varying quantities of questions, up to a maximum of 100. The results are presented in Figure. 3. We do not present results for in-context learning methods due to their poor performance.
In terms of precision, baseline methods show a significant decrease as the number of generated questions increases, indicating a deterioration in semantic consistency. In contrast, the precision of the proposed methods remains consistently high, with only a minor decrease observed when generating 100 questions. This highlights the superior semantic consistency of the proposed method, which is capable of generating a larger set of questions. For recall, our proposed methods achieve a significant improvement, increasing from $0.82$ to above $0.89$ as more questions are generated. Notably, our methods reach a recall of approximately $0.82$ when generating just 10 questions, surpassing baseline methods when generating 100 questions. This demonstrates that the “one-to-many” training objective effectively enables LLM to explore the semantic space surrounding the source question.
@@
Algorithm Greedy Algorithm for Maximizing Pairwise Diversity
2:Subset $S\subseteq Q^{*}$
[…]
5:     Select $q^{*}\in Q^{*}\setminus S$ that maximizes:
[…]
7:         Add $q^{*}$ to $S$
[…]
9:     end if
10:end while
11:return $S$
@@
Table 3: Comparison of random selection and greedy algorithm in efficiency and semantic diversity.
@@ 4.3 Selection of Generated Questions
Building on the effectiveness of the proposed SQG methods, we now delve into the selection of generated questions for knowledge base augmentation. Instead of generating the targeted number of questions directly, we argue that a two-step approach, which generates many candidate questions and then selects the best samples, would offer greater flexibility and achieve better results in industrial applications. With the detailed problem formulation and proof of submodularity and NP-hardness presented in Appendix E, we propose a greedy algorithm that efficiently approximates the optimal solution with a guaranteed approximation bound of $1-1/e$. The greedy algorithm iteratively selects the sample $q^{*}$ that provides the highest marginal gain relative to its cost while satisfying the budget constraint.
The comparison of random and greedy selection methods is presented in Table 3, which highlights the superior diversity compared to random sampling. Given the augmentation is an offline process, we argue that the time invested in greedy search is a worthwhile trade-off for the improvement in total diversity. We also note that the brute-force method, which evaluates all possible subsets and guarantees optimality, is inefficient for subset selection due to the NP-hardness of the problem, where selecting 5 questions from a set of 20 generated questions requires evaluating 15,504 combinations.
@@ 5 Conclusion
In this work, we present a comprehensive approach to expanding the knowledge base of service chatbots by generating similar questions using Large Language Models. By introducing a one-to-many training objective, we enabled the effective exploration of the semantic space surrounding the source question. Additionally, by utilizing customer intention as contextual guidance, the generation of similar questions presented greater semantic diversity while still adhering to the customer’s intention. Both quantitative and qualitative experiments on the production dataset demonstrate the superior performance of our proposed methods. Furthermore, the optimization framework enables our method to be seamlessly integrated into existing production systems, offering enhanced flexibility and computational efficiency. Future research is encouraged to apply the SQG methods to the Retrieval Augmented Generation (RAG) knowledge base for improved customer interactions with LLMs.
@@ Appendix B Implementation Details
We utilize the lightweight ChatGLM2-6B Zeng et al. (2023) as our base model for its superior performance in the Chinese language. Instead of employing parameter-efficient tuning Ding et al. (2023), we implemented full tuning using 8 $\times$ Nvidia A100 GPUs on 90,000 training instances, as this approach yields significantly better performance. In the case of SimBERT and RoFormer-Sim, we set the temperature and top-k parameters to 0.9 and 5, respectively. For the ChatGLM2-based methods, we adhere to the default generation settings. For our proposed methods, the number of generated outputs, denoted as $L$, is set to 20, as we have observed that increasing this number often leads to degraded performance, which is discussed in Section 4.2.
{CJK}
UTF8gbsn
Method High precision Low precision
SimBERT 1. 证明要多久才可以开? (How long does take to obtain a certificate?) 2. 开证明一般要多久才能拿到? (How long does it generally take to obtain a certificate?) 3. 一般证明需要多久才可以开? (How long does it generally take to issue a certificate?) 1. 证明要多长时间?(How long does it take for the proof?) 2. 公司证明怎么开?(How do I go about obtaining proof from the company?) 3. 证明书需要几个证明时间? (How many proofing sessions are required for the proofing book?)
Appendix C Qualitative Example Demonstration
We demonstrate a typical example from customer service support to further illustrate the effectiveness of the proposed batch-generation methods, as shown in Table 4. We provide three questions generated with high precision and three with low precision relative to the source question. The high-precision examples across all methods demonstrate strong semantic consistency with the source question, while the low-precision examples reveal notable distinctions. Questions generated by SimBERT deviate from the source question and exhibit a lack of fluency, exemplified by ‘How do I go about obtaining proof from the company?’ ( {CJK}UTF8gbsn‘公司证明怎么开?’). This departure from the original question is consistent with the significant drop in the recall score as seen earlier in Figure. 3. Conversely, the Context-Aware Batch Eeneration method can generate novel expressions, such as ‘I applied for the certificate, why haven’t I received it yet?’ ( {CJK}UTF8gbsn‘我开了证明怎么还没收到?’), which suggests a more effective exploration of the semantic space surrounding the source question. In the case of the Intention-Enhanced Batch Generation method, it becomes evident that information derived from the intention, not present in the source question, is effectively harnessed to generate similar questions, as shown by ‘electronic certificate’ ( {CJK}UTF8gbsn‘电子证明’) and ‘paper certificate’ ({CJK}UTF8gbsn‘纸质证明’). This underscores the importance of the intention as an effective way to enhance the diversity of generated questions.
@@ Appendix C Qualitative Example Demonstration
Appendix D More Results
In this section, we present additional experimental results based on domain-specific datasets to gain further insights into industrial deployment. The objective of these experiments is to demonstrate two crucial properties of the proposed methods: 1) Model invariance, which ensures that the performance of the proposed methods remains consistent regardless of the model type or size applied, and 2) Domain invariance, which ensures that the performance of the proposed methods remains consistent across different domains or tasks. To explore these aspects, we incorporate two additional LLMs through OpenAI API, GPT-3.5 and GPT-4, which are more advanced models with enhanced capabilities. Given that these models are black-boxed, we implement only the few-shot setting while incorporating batch generation to reinforce the context. We utilize three publicly available QA datasets covering distinct domains: telecommunications (telecom), banking services for loans (loan), and legal services (legal). For each domain, we generate 20 similar questions for each source question [^2] Sample code and datasets can be found at https://anonymous.4open.science/status/SQGwithLLMs.
The results presented in Table 5 demonstrate the varying performances of different models. GPT-4 generally outperforms GPT-3.5 due to its larger model size. However, ChatGLM2 often exhibits competitive performance, especially with intent enhancement, attributed to task-specific fine-tuning. This highlights the importance of domain adaptation in achieving better text quality. However, larger LLMs show advantages when evaluated on generic domains, such as legal and law, benefiting from extensive pretraining data.
The generation speed of the methods varies considerably. Both proposed methods benefit from the proposed one-to-many batch generation, resulting in significantly faster generation compared to the one-to-one approach. For instance, the average speed for Context-Based Batch Generation using GPT-3.5 is 5.83 seconds per item (i.e., source question), while Intention-Enhanced Batch Generation takes 6.35 seconds. In contrast, the same task using the one-to-one paradigm requires 50.23 seconds per item, significantly slower than the proposed methods. We do not present results for one-to-one methods due to their poor performance and slow generation.
Finally, we observe that the relative performance in terms of generation quality, speed, and retrieval capability remains consistent across different tasks and models. The larger models tend to produce better results, which is consistent with previous research; however, the improvements are not significantly large. Due to cost constraints, we still recommend a smaller model for accomplishing the Similar Question Generation tasks.
Appendix E Optimization Framework with Resource Constraints
@@
E.1 Problem Formulation
The optimization framework is designed to maximize semantic diversity within a constrained resource budget $B$. Here, $B$ serves as a critical parameter representing the system’s capacity for handling selected questions, which could be interpreted as either:
  • Storage Constraint: The total storage capacity available to maintain the selected questions, where the length of each question (e.g., number of characters) contributes to its storage cost.
  • Retrieval Power Constraint: The computational resources or time available for retrieving answers. The more questions selected, the higher the retrieval complexity, as the system must compute pairwise distances or other similarity measures.
@@ E.1 Problem Formulation
[…]
E.2 Practical Implications and Advantages of Budget Constraints
@@
Customizability and Flexibility
The budget acts as a tunable parameter that can be adapted to specific application requirements. By adjusting $B$, practitioners can fine-tune the trade-off between diversity, storage, and retrieval efficiency:
  • Trade-off Optimization: The ability to customize $B$ allows decision-makers to align system performance with operational priorities, whether optimizing for computational speed, minimizing costs, or maximizing diversity.
  • Flexible Deployment: The framework’s design supports diverse system configurations, making it suitable for applications with varying capacities, from resource-constrained environments to large-scale systems.
In summary, the introduction of $B$ not only addresses practical constraints in real-world applications but also provides a flexible and scalable solution for optimizing diversity under limited resources. The framework’s adaptability ensures its relevance across a wide range of use cases, from resource-limited systems to high-performance environments.
E.3 Proposed Solution
In this subsection, we first establish that the problem is NP-hard by reducing it from a well-known NP-hard problem. Next, we prove that the objective function is submodular, which allows us to design an efficient greedy algorithm. Finally, we demonstrate the $1-1/e$ approximation bound of the proposed Greedy Algorithm in Section 4.3.
E.3.1 NP-Hardness of the Problem
Theorem E.1.
@@ E.3.1 NP-Hardness of the Problem
We establish the NP-hardness of the problem by reducing it from the Maximum Diversity Problem (MDP), which is a known NP-hard problem Kuo et al. (1993). The MDP is defined as selecting $k$ elements from a set such that the sum of pairwise distances is maximized:
@@
In our problem, consider the special case where the cost of each element is uniform, i.e., $\text{cost}(q)=1$ for all $q\in Q^{*}$, and the budget constraint is $B=k$. Under these conditions, the budget constraint simplifies to selecting exactly $k$ elements from $Q^{*}$. This special case is equivalent to the MDP. Since the MDP is NP-hard, and our problem generalizes it by introducing arbitrary costs and a more flexible budget constraint, our problem is at least as hard as the MDP.
Moreover, the problem is in NP because, given a candidate solution $S$, it is straightforward to:
  • Verify in polynomial time whether $\sum_{q\in S}\text{cost}(q)\leq B$, and
  • Compute the total diversity $\sum_{\begin{subarray}{c}q_{a},q_{b}\in S\\ q_{a}\neq q_{b}\end{subarray}}\text{dist}(q_{a},q_{b})$.
Therefore, the problem is NP-hard. ∎
E.3.2 Submodularity of the Objective Function
Although the problem is NP-hard, we observe that the objective function is submodular, a property that enables the use of an efficient greedy algorithm to approximate the optimal solution.
Theorem E.2.
@@ E.3.2 Submodularity of the Objective Function
Let $A\subseteq B\subseteq Q^{*}$. Then, for any $x\notin B$:
$\Delta f(A,x)=\sum_{q\in A}\text{dist}(x,q),$
$\quad\Delta f(B,x)=\sum_{q\in B}\text{dist}(x,q).$
Since $A\subseteq B$, the terms in $\Delta f(A,x)$ are a subset of those in $\Delta f(B,x)$. Therefore, the marginal gain from adding $x$ diminishes as the set grows, satisfying the submodularity condition:
@@
Thus, $f(S)$ is submodular. Moreover, $f(S)$ is non-decreasing, as adding a new element can only increase the sum of pairwise distances. ∎
E.4 Complexity and Approximation Analysis
The proposed greedy algorithm in Section 4.3 efficiently selects a diverse subset under a budget constraint while achieving a provable approximation guarantee.
E.4.1 Approximation Guarantee

### v3

@@ 6.1 Main Results
Table 3 presents results for generating 20 similar questions. Most methods achieve high precision, with generated questions closely aligning with source semantics. However, baseline methods show low recall, indicating limited diversity as a key challenge in SQG. Static in-context learning methods underperform in both precision and recall due to irrelevant demonstrations. Fine-tuning with a one-to-one objective (ChatGLM2-FT) improves recall while maintaining precision, demonstrating the value of task-specific adaptation. The proposed one-to-many training objective (Context-Aware) enhances both precision and recall, and the Intention-Enhanced method further improves diversity and relevance. The inclusion of dynamic demonstration selection achieves state-of-the-art performance, surpassing zero-shot methods.
Human evaluation shows that general-purpose text-generation models, SimBERT and RoFormer-Sim, perform poorly, with only 20% of generated questions meeting the acceptance criteria. ChatGLM2-FT improves this to 37.9%, but still remains largely redundant and fails to meet practical needs. While the Context-Aware method excels in quantitative metrics, its impact on the acceptance ratio is modest. Introducing the customer’s intention via the source answer significantly broadens the space of exploration, resulting in the largest number of usable similar questions. This emphasizes the importance of contextual information in improving relevance and diversity.
Finally, character-level diversity shows that ChatGLM2-FT and the proposed methods outperform SimBERT, RoFormer-Sim, and ChatGLM2 (Few-Shot) in Distinct scores. Although the zero-shot ChatGLM2 achieves the highest Distinct score, it sacrifices consistency, as reflected in its low precision score due to the unconditioned generation process, which is not an ideal behavior.
Figure 3: Performance comparison of similar question generation methods with varying number of questions.
@@
In real applications, the desired quantity of similar questions often falls within the range of several tens to hundreds. Therefore, we systematically examine the performance of the proposed method by generating varying quantities of questions, up to a maximum of 100.
As shown in Figure 3, the Intention-Enhanced approach shows a surprising trend: precision stays consistently high, with only a slight decrease when generating up to 100 questions, while baseline methods experience a significant precision drop as the number of questions grows. This strength comes from the approach’s ability to balance relevance and variety, creating a diverse set of questions that closely match user intents and cover a broad range of query expressions. In contrast, baseline methods lose semantic consistency as question volume increases, reducing relevance and lowering precision. Additionally, our approach greatly improves recall, rising from $0.82$ to above $0.89$ as more questions are generated. Notably, our methods reach a recall of approximately $0.82$ when generating just 10 questions, surpassing baseline methods when generating 100 questions. This demonstrates that the one-to-many training objective effectively enables LLM to explore the semantic space surrounding the source question.
@@
Table 4: Comparison of random selection and greedy algorithm in time efficiency and semantic diversity.
Method High precision Low precision
SimBERT 1. 证明要多久才可以开? (How long does it take to obtain a certificate?) 2. 开证明一般要多久才能拿到? (How long does it generally take to obtain a certificate?) 3. 一般证明需要多久才可以开? (How long does it generally take to issue a certificate?) 1. 证明要多长时间?(How long does it take for the proof?) 2. 公司证明怎么开?(How do I go about obtaining proof from the company?) 3. 证明书需要几个证明时间? (How many proofing sessions are required for the proofing book?)
@@ 6.3 Selection of Generated Questions
To select a subset of similar questions for knowledge base expansion, comparison in Table 4 highlights the superior diversity of the proposed greedy selection algorithm. Given that the augmentation is an offline process, we argue that the time invested in greedy search is a worthwhile trade-off for the improvement in total diversity. We also note that the exhaustive search method, which evaluates all possible subsets and guarantees optimality, is inefficient due to the NP-hardness of the problem, where selecting five questions from a set of 20 generated questions requires evaluating 15,504 combinations.
6.4 Application Performance
We evaluated the performance of our system in realistic customer service applications within the banking industry, deploying it for an in-app chatbot serving over 100,000 active users over a three-month period. The chatbot, equipped with a knowledge base augmented with the Intention-Enhanced approach, achieved the highest service success rate, with over 92% user satisfaction, outperforming the unaugmented knowledge base at 74% and the simple context-aware approach at 83.83%. These results reveal the effectiveness of proposed augmentation strategies in enhancing service quality through improved query matching accuracy and more semantically consistent response retrieval.
7 Qualitative Example Demonstration
We demonstrate a typical example from customer service support to illustrate the effectiveness of the proposed methods, as shown in Table 5. The high-precision examples across all methods demonstrate strong semantic consistency with the source question, while the low-precision examples reveal notable distinctions. Questions generated by SimBERT deviate from the source question and exhibit a lack of fluency, exemplified by ‘How can I obtain a certificate from the company?’ ( ‘公司证明怎么开?’). This departure from the original question is consistent with the significant drop in the recall score as seen earlier in Figure 3. Conversely, the Context-Aware Batch Generation method can generate novel expressions, such as ‘I applied for the certificate, why haven’t I received it yet?’ ( ‘我开了证明怎么还没收到?’), which suggests a more effective exploration of the semantic space surrounding the source question. In the case of the Intention-Enhanced Batch Generation method, it becomes evident that information derived from the intention, not present in the source question, is effectively harnessed to generate similar questions, as shown by ‘electronic certificate’ ( ‘电子证明’) and ‘paper certificate’ (‘纸质证明’). This underscores the importance of the customer intention as an effective guide to enhance the diversity of generated questions.
@@ 8 Conclusion
This work presents an innovative approach to expanding the knowledge base of compliance-guaranteed service chatbots by generating similar questions with LLMs. By introducing a one-to-many training objective and utilizing customer intention as contextual guidance, we enhanced semantic diversity while staying aligned with the customer’s intent. The optimization framework enables our method to be seamlessly integrated into existing production systems, offering great flexibility and efficiency. These promising findings highlight LLMs’ growing role in augmenting conventional system architectures in scenarios where standalone LLMs are not directly applicable, encouraging further research into LLM-guided systems for industrial deployment.
Limitations
While this work pioneers a novel strategy for augmenting a retrieval-based chatbot system with LLM-generated similar questions, it has two main limitations. First, it assumes customer inquiries are monolingual, overlooking the challenges of multilingual query matching, increasingly common in multinational enterprises and domains with frequent code-switching Jiang et al. (2016). Second, human evaluations by domain experts are costly and lack scalability. Recent studies suggest using LLM-as-a-judge to replace human involvement in performance evaluation, which could be better integrated into the proposed method to provide human-aligned feedback during question generation.
Acknowledgments
This paper is partially supported by several ongoing projects led or coordinated by Prof. Zhang Chen, including P0048887 (Innovation and Technology Fund - ITSP, ITS/028/22FP), P0051906 (RGC Early Career Scheme, 25600624), and P0054482 (Two Square Capital Limited donation).
@@ Appendix B Implementation Details
We utilize the lightweight ChatGLM2-6B Zeng et al. (2023) as our base model for its superior performance in the Chinese language Hong et al. (2025b). Instead of employing parameter-efficient tuning Ding et al. (2023), we implemented full tuning using Nvidia A100 GPUs on 90,000 training instances, as this approach yields significantly better performance. In the case of SimBERT and RoFormer-Sim, we set the temperature and top-k parameters to 0.9 and 5, respectively. For the ChatGLM2-based methods, we adhere to the default generation settings. The number of generated outputs, denoted as $L$, is set to 20, as we have observed that increasing this number often leads to degraded performance, which is discussed in Section 6.2.
@@
Appendix C More Results
In this section, we present additional experimental results to gain further insights into industrial deployment and the general applicability of the proposed methods. The objective of these experiments is to demonstrate two crucial properties: 1) Model invariance, which ensures that the performance remains consistent regardless of the model type or size applied, and 2) Domain invariance, which ensures that the performance remains consistent across different domains or tasks.
To explore these aspects, we incorporate two additional LLMs through the OpenAI API, GPT-3.5 and GPT-4, which are proprietary models with advanced capabilities. Given that these models are black-boxed, we implement the Intention-Enhanced prompting without model finetuning. For evaluation, we utilize three publicly available QA datasets covering distinct domains: telecommunications (telecom), banking services for loans (loan), and legal services (legal). For each source question, we generate 20 similar questions.
The results presented in Table 6 demonstrate the varying performances of different models. GPT-4 generally outperforms GPT-3.5 due to its larger model size. However, the fine-tuned ChatGLM2 often exhibits competitive performance, especially with intent enhancement, attributed to task-specific fine-tuning. This highlights the importance of domain adaptation in achieving better text quality. However, larger LLMs show advantages when evaluated on generic domains, such as legal and law, benefiting from extensive pretraining data.
The generation speed of the methods varies considerably. Both proposed methods benefit from the one-to-many batch generation, resulting in significantly faster generation compared to the one-to-one approach. For instance, the average speed for Context-Based Batch Generation using GPT-3.5 is 5.83 seconds per item (i.e., source question), while Intention-Enhanced Batch Generation takes 6.35 seconds. In contrast, the same task using the one-to-one paradigm requires 50.23 seconds per item, significantly slower than the proposed methods. We do not present results for one-to-one methods due to their poor performance and slow generation.
Finally, we observe that the relative performance in terms of generation quality, speed, and retrieval capability remains consistent across different tasks and models. The larger models tend to produce better results, which is consistent with previous research; however, the improvements are not significantly large. Due to cost constraints, we still recommend a smaller fine-tuned model for accomplishing the Similar Question Generation tasks.
Appendix D Similar Question Selection Optimization
@@
D.1 Problem Formulation
The optimization framework is designed to maximize semantic diversity within a constrained resource budget $B$, a key parameter representing the system’s capacity to manage selected questions. This budget can be interpreted as either a storage constraint, limiting the number of questions based on their length (e.g., number of characters), or a retrieval power constraint, where computational resources or time available for retrieving answers are limited. As more similar questions are selected, both storage and retrieval complexity are elevated, leading to high maintenance costs and increased system latency due to the need to compute similarity measures.
@@ D.1 Problem Formulation
$\text{s.t.}\quad\sum_{q\in S}\text{cost}(q)\leq B.$
D.2 Practical Implications and Advantages of Budget Constraints
@@
Flexibility
The budget acts as a tunable parameter that can be adapted to specific application requirements. By adjusting $B$, practitioners can fine-tune the trade-off between diversity, storage, and retrieval efficiency.
D.3 Proof of Proposed Solution
We first establish that the problem is NP-hard by reducing it from a well-known NP-hard problem. Next, we prove that the objective function is submodular, enabling the use of the proposed greedy algorithm as described in Algorithm 2. Finally, we demonstrate the 11/e1-1/e approximation bound of the Greedy Algorithm.
D.3.1 NP-Hardness of the Problem
Theorem D.1.
@@ D.3.1 NP-Hardness of the Problem
We establish NP-hardness by reducing the problem from the Maximum Diversity Problem (MDP), a known NP-hard problem Kuo et al. (1993). The MDP involves selecting $k$ elements from a set to maximize the sum of pairwise distances:
@@
In our problem, consider the special case where each element has uniform cost, i.e., $\text{cost}(q)=1$ for all $q\in Q^{*}$, and the budget constraint is $B=k$. This simplifies to selecting exactly $k$ elements from $Q^{*}$, equivalent to the MDP. Since the MDP is NP-hard, and our problem generalizes it with arbitrary costs and a flexible budget constraint, our problem is at least as hard as the MDP. Moreover, the problem is in NP, as verifying a candidate solution $S$ involves checking in polynomial time whether $\sum_{q\in S}\text{cost}(q)\leq B$ and computing the total diversity $\sum_{\begin{subarray}{c}q_{a},q_{b}\in S\\ q_{a}\neq q_{b}\end{subarray}}\text{dist}(q_{a},q_{b})$. Thus, the problem is NP-hard. ∎
D.3.2 Submodularity of the Objective Function
Although the problem is NP-hard, the objective function is submodular, a property that enables an efficient greedy algorithm to approximate the optimal solution.
Theorem D.2.
@@ D.3.2 Submodularity of the Objective Function
For $A\subseteq B\subseteq Q^{*}$ and $x\notin B$:
$\Delta f(A,x)=\sum_{q\in A}\text{dist}(x,q)$
$\Delta f(B,x)=\sum_{q\in B}\text{dist}(x,q).$
Since $A\subseteq B$, the terms in $\Delta f(A,x)$ are a subset of those in $\Delta f(B,x)$. Thus, the marginal gain from adding $x$ decreases as the set grows, satisfying the submodularity condition:
@@
Hence, $f(S)$ is submodular. Additionally, $f(S)$ is non-decreasing, as adding an element can only increase the sum of pairwise distances. ∎
D.4 Complexity and Approximation Analysis
The proposed greedy algorithm in Algorithm 2 efficiently selects a diverse subset under a budget constraint while achieving a provable approximation guarantee.
D.4.1 Approximation Guarantee